# 2.5D Transformer: An Efficient 3D Seismic Interpolation Method without Full 3D Training

### v1

@@ 2.1 Transformer
As a classical DL framework, Transformer has gained significant attention from academia and industry in recent years. It has been applied to the fields of natural language processing [vaswani2017attention] and computer vision [dosovitskiy2020image]. Owing to the core operation, MSA, Transformer is capable of capturing the global contextual information and shows better performance than CNNs in numerous cases.
@@ 2.2.1 T-2D/3D
In Fig. 2a, the T-2D is composed of two 3×3 convolutional layers (Convs), a head block (HB), four 2D Transformer encoders, and a tail block (TB). A 2D patch with a size of 40×40 is firstly input into the first Conv, and the channel number is increased from 1 to 32. Then, the features with a size of 40×40×32 are input into HB whose structure is shown in Figure 2b. The HB consists of two identical parts, and each one comprises two Convs, a rectified linear unit (ReLU), and a residual connection. The output features of HB are propagated into the core module of T-2D: four successively-connected Transformer encoders, thereby capturing long-range dependencies. Moreover, dense connections [Huang_2017_CVPR], whose effectiveness has been validated by [dong2025deep], are deployed to enrich the feature interactions among the four Transformer encoders. Subsequently, the TB displayed in Fig. 2c is used for final feature refinement, which is composed of three Convs interleaved with leaky ReLUs in a feed-forward manner. The TB enables efficient non-linear transformation while mitigating the risk of neuron inactivation. The output features of TB and HB are fused via a residual connection, thus avoiding the phenomenon of gradient vanishing. Finally, to generate the final output feature, the last Conv is used to integrate features across channel dimension and reduce the channel number back to one.
@@
The T-2.5D is generated by adding four 3D SDAs to T-2D. In Fig. 4a, we place a 3D SDA before each 2D Transformer encoder. As shown in Fig. 4b, the 3D SDA is composed of two linear layers and a depth-wise 3D Conv (DWConv3d). Specifically, the first linear layer is used to project the input feature into a 3D space, allowing the following DWConv3d to capture abundant spatial contextual information. At the end, another linear layer projects the feature back to the original dimension.
@@ 2.4 Transformer-based Interpolation Theory
where and denote the weights and biases, respectively, is the batch size (i.e., the pair number of training data patches used in one iteration), and represent the -th paired training data patches, and denotes the L2 norm. In this paper, we use the ADAM algorithm [kingma2014adam] to update the trainable parameters in an iterative manner. Upon obtaining the optimal parameters , we can generate the interpolated result by inputting the incomplete data into the nonlinear relationship :
@@ 3.2 Metrics for Interpolation Results
In this study, the peak signal-to-noise ratio (PSNR) and structural similarity index measure [SSIM; wang2004image] are used to evaluate the performance of different methods quantitatively. The PSNR of 2D data is defined as the following equation:
@@
SSIM is a commonly used metric to evaluate the perceptual similarity between two data. It considers the changes in structural information, luminance, and contrast [wang2004image]. Given a 2D/3D complete data and its corresponding interpolated result , SSIM is defined as:
@@ 3.3.2 Testing results
As shown in Fig. 7a, the test data is of shape 224×224×20 extracted from Fig. 6a. We randomly delete 50% of the traces to generate the incomplete data in Fig. 7e. Fig. 7b-7d presents the reconstruction results. Fig. 7f–7h display the residual images between the interpolated results (Fig. 7b–7d) and the complete data (Fig. 7a), and Fig. 7i–7k are the local similarity [fomel2007local] maps between Fig. 7f–7h and Fig. 7a. Both of them are utilized to measure the signal leakage of the interpolated results, and lower energy indicates better interpolation performance. We can observe that all the three methods are able to reconstruct the missing traces to some extent. However, T-2D exhibits stronger signal leakage than T-3D and T-2.5D as shown in Fig. 7f and 7i. In comparison, T-2.5D achieves a comparable performance to T-3D, which validates the effectiveness of T-2.5D.
For a clearer comparison, the sixth line is plotted in 2D (Fig. 8). Due to the lack of spatial correlation information, T-2D yields the poorest performance. This is evidenced by its significantly lower quantitative results and the most severe signal leakage among all networks, as shown in Fig. 8c and 8d. In comparison, T-2.5D achieves an interpolation result close to that of T-3D, and both of their residual images show visually little signal leakage.
Furthermore, we plot in Fig. 9 the f-k spectra of the complete data (Fig. 8a), 50% randomly sampled data (Fig. 8b), the three interpolated results (Fig. 8c, 8e, and 8g), and the corresponding residual images (Fig. 8d, 8f, and 8h). As illustrated in Fig. 9c–9e, the aliased energy presented in corrupted data (Fig. 9b) has been largely removed by all the three methods. However, as shown in Fig. 9c, stronger residual interference exists in the result of T-2D. It is also observed in Fig. 9h that the energy leakage of T-2.5D is evidently weaker than that of T-2D, and is very close to that of T-3D.
@@ 3.4.2 Testing results
Similarly, the tenth line is illustrated in Fig. 12 as a 2D view. The numerical results in Fig. 12g and little leakage in Fig. 12h indicate the great interpolation ability of T-2.5D.
Fig. 13 displays the f-k spectrum of the complete data (Fig. 12a), 50% randomly sampled data (Fig. 12b), the interpolated results in Fig. 12c, 12e and 12g, respectively, and the corresponding residual images (Fig. 12d, 12f, and 12h). T-2D still produces more residual interference than the other two methods as indicated by the red arrows and ovals in Fig. 13c. In comparison, the recovered spectrum in Fig. 13e and the minimal energy leakage in Fig. 13h demonstrate the strong interpolation capability of T-2.5D.
@@ 3.5 Analysis of the computational cost
Generalization is an important ability that allows a DL model to effectively process variant datasets that are not used during training [zhang2021understanding]. To further explore the generalization capability, the Opunake dataset, which is from another 3D survey in New Zealand, is utilized for further TL of T-2.5D and T-3D. Specifically, we extract approximately 1000 and 250 volumes with a shape of 40×40×16 from the full Opunake dataset for training and validation, respectively, and extract a volume with a shape of 224×224×20 as test data. Here, TL for T-2.5D and T-3D is performed in a full fine-tuning manner. The comparisons of the interpolated results before and after the TL are presented in Fig. 14. As indicated by the red boxes in Fig. 14c and 14d, the interpolated results before applying TL contain certain discontinuous events. As illustrated in Fig. 14e and 14f, this issue is alleviated after applying TL. The numerical results have also been evidently improved. Meanwhile, Fig. 15 gives the corresponding residual images and local similarity maps. Compared to Fig. 15a–15d, the lower energy in Fig. 15e–15h indicates better interpolation performance after applying TL.
We plot the 20th line of the test data from Fig. 14 for 2D comparison. As shown in Fig. 16c and 16e, the two interpolated results before applying TL contain discontinuous events as indicated by the red boxes. After applying TL, the continuity and smoothness have been improved, as shown in Fig. 16g and 16i. T-2.5D still achieves comparable performance to that of T-3D, with an even higher SSIM value. In summary, T-2.5D has great potential for effectively interpolating randomly missing traces in 3D seismic data from different survey areas, exhibiting great generalization ability.
@@ 4.1 Testing in the Regularly Missing Case
All the interpolated results in the previous section are dealing with irregularly missing traces. In fact, T-2.5D can also be leveraged to handle regular gaps to increase the acquisition density. To validate this, we further assesse the proposed T-2.5D on the Kerry dataset with 67% regularly missing traces (removing two out of every three traces) in Fig. 17b. As shown in Fig. 17c, 17e, and 17g, the interpolation result of T-2.5D shows superior performance to T-2D. Furthermore, it has achieved even better numerical results than T-3D does under such a high missing rate. The residual image in Fig. 17h further demonstrates that T-2.5D can effectively recover signals with little leakage in the regularly missing scenario.
@@ 4.2 Comparison of Trainable Parameters
The number of trainable parameters in a network model is also an important factor for DL-based methods. Models with more trainable parameters can represent more complex functions, but may require more memory and computation [tu2024overview]. We further compare the trainable parameters of the three networks. Notably, the total trainable parameters of T-2.5D are composed of two parts: the ones from the four Transformer encoders in the first stage, and the ones from the remainder of T-2.5D in the second stage. As shown in Table 5, the trainable parameters surge from 2D to 3D. Although T-2D has the fewest trainable parameters, its interpolation performance is also inferior to the other two networks. Although the total trainable parameters of T-2.5D are only 75.34% of those of T-3D, T-2.5D can achieve performance comparable to T-3D, or even surpassing it in certain cases. This finding further supports the lightweight and efficient nature of T-2.5D.

### v2

@@ 2.1 Transformer
As a classical DL framework, Transformer has gained significant attention from academia and industry in recent years. It has been applied to the fields of natural language processing (Vaswani et al., 2017) and computer vision (Dosovitskiy, 2020). Owing to the core operation, MSA, Transformer is capable of capturing the global contextual information and shows better performance than CNNs in numerous cases.
@@ 2.2.1 T-2D/3D
In Fig. 2a, the T-2D is composed of three 3×3 convolutional layers (Convs), a head block (HB), four 2D Transformer encoders, and a tail block (TB). A 2D patch with a size of 40×40 is firstly input into the first Conv, and the channel number is increased from 1 to 32. Then, the features with a size of 40×40×32 are input into HB whose structure is shown in Figure 2b. The HB consists of two identical parts, and each one comprises two Convs, a rectified linear unit (ReLU), and a residual connection. The output features of HB are propagated into the core module of T-2D: four successively-connected Transformer encoders, thereby capturing long-range dependencies. Moreover, dense connections Huang et al. (2017), whose effectiveness has been validated by Dong et al. (2025), are deployed to enrich the feature interactions among the four Transformer encoders. Then, a Conv is used to refine the global feature, and an LN layer is deployed to stabilize the training. Subsequently, the TB displayed in Fig. 2c is used for final feature refinement, which is composed of three Convs interleaved with leaky ReLUs in a feed-forward manner. The TB enables efficient non-linear transformation while mitigating the risk of neuron inactivation. The output features of TB and HB are fused via a residual connection, thus avoiding the phenomenon of gradient vanishing. Finally, to generate the final output feature, the last Conv is used to integrate features across channel dimension and reduce the channel number back to one.
@@
The T-2.5D is generated by adding four 3D SDAs to T-2D. In Fig. 4a, we place a 3D SDA before each 2D Transformer encoder. As shown in Fig. 4b, the 3D SDA is composed of two linear layers and a depth-wise 3D Conv (DWConv3d). Specifically, the first linear layer is used to project the input feature into a 3D space, allowing the following DWConv3d to capture abundant spatial contextual information. Another linear layer is used to project the feature back to the original dimension. A residual connection is utilized to send the input to the output, to prevent potential information loss.
@@ 2.4 Transformer-based Interpolation Theory
where and denote the weights and biases, respectively, is the batch size (i.e., the pair number of training data patches used in one iteration), and represent the -th paired training data patches, and denotes the L2 norm. In this paper, we use the ADAM algorithm Kingma and Ba (2014) to update the trainable parameters in an iterative manner. Upon obtaining the optimal parameters , we can generate the interpolated result by inputting the incomplete data into the nonlinear relationship :
@@ 3.2 Metrics for Interpolation Results
In this study, the peak signal-to-noise ratio (PSNR) and structural similarity index measure (SSIM; Wang et al., 2004) are used to evaluate the performance of different methods quantitatively. The PSNR of 2D data is defined as the following equation:
@@
SSIM is a commonly used metric to evaluate the perceptual similarity between two data. It considers the changes in structural information, luminance, and contrast (Wang et al., 2004). Given a 2D/3D complete data and its corresponding interpolated result , SSIM is defined as:
@@ 3.3.2 Testing results
As shown in Fig. 7a, the test data is of shape 224×224×20 extracted from Fig. 6a. We randomly delete 50% of the traces to generate the incomplete data in Fig. 7e. Fig. 7b-7d presents the reconstruction results. Fig. 7f–7h display the residual images between the interpolated results (Fig. 7b–7d) and the complete data (Fig. 7a), and Fig. 7i–7k are the local similarity (Fomel, 2007) maps between Fig. 7f–7h and Fig. 7a. Both of them are utilized to measure the signal leakage of the interpolated results, and lower energy indicates better interpolation performance. We can observe that all the three methods are able to reconstruct the missing traces to some extent. However, T-2D exhibits stronger signal leakage than T-3D and T-2.5D as shown in Fig. 7f and 7i. In comparison, T-2.5D achieves a comparable performance to T-3D, which validates the effectiveness of T-2.5D.
For a clearer comparison, the sixth line is plotted in 2D (Fig. 8). Due to the lack of spatial correlation information, T-2D yields the poorest performance. This is evidenced by its significantly lower quantitative results and the most severe signal leakage among all networks, as shown in Fig. 8b and 8f. In comparison, T-2.5D achieves an interpolation result close to that of T-3D, and both of their residual images show visually little signal leakage.
Furthermore, we plot in Fig. 9 the f-k spectra of the complete data (Fig. 8a), 50% randomly sampled data (Fig. 8e), the three interpolated results (Fig. 8b–8d), and the corresponding residual images (Fig. 8f–8h). As illustrated in Fig. 9b–9d, the aliased energy presented in corrupted data (Fig. 9e) has been largely removed by all the three methods. However, as shown in Fig. 9f, stronger residual interference exists in the result of T-2D. It is also observed in Fig. 9h that the energy leakage of T-2.5D is evidently weaker than that of T-2D, and is very close to that of T-3D.
@@ 3.4.2 Testing results
Similarly, the tenth line is illustrated in Fig. 12 as a 2D view. The numerical results in Fig. 12d and little leakage in Fig. 12h indicate the great interpolation ability of T-2.5D.
Fig. 13 displays the f-k spectrum of the complete data (Fig. 12a), 50% randomly sampled data (Fig. 12e), the interpolated results in Fig. 12b–12d, respectively, and the corresponding residual images (Fig. 12f–12h). T-2D still produces more residual interference than the other two methods as indicated by the red arrows and ovals in Fig. 13b. In comparison, the recovered spectrum in Fig. 13d and the energy leakage in Fig. 13h demonstrate the strong interpolation capability of T-2.5D.
@@ 3.5 Analysis of the computational cost
Generalization is an important ability that allows a DL model to effectively process variant datasets that are not used during training (Zhang et al., 2021). To further explore the generalization capability, the Opunake dataset, which is from another 3D survey in New Zealand, is utilized for further TL of T-2.5D and T-3D. Specifically, we extract approximately 1000 and 250 volumes with a shape of 40×40×16 from the full Opunake dataset for training and validation, respectively, and extract a volume with a shape of 224×224×20 as test data. Here, TL for T-2.5D and T-3D is performed in a full fine-tuning manner. The comparisons of the interpolated results before and after the TL are presented in Fig. 14. As indicated by the red boxes in Fig. 14b and 14e, the interpolated results before applying TL contain certain discontinuous events. As illustrated in Fig. 14c and 14f, this issue is alleviated after applying TL. The numerical results have also been evidently improved. Meanwhile, Fig. 15 gives the corresponding residual images and local similarity maps. Compared to Fig. 15a, 15e, 15b, 15f, the lower energy in Fig. 15c, 15g, 15d, 15h indicates better interpolation performance after applying TL.
We plot the 20th line of the test data from Fig. 14 for 2D comparison. As shown in Fig. 16b and 16c, the two interpolated results before applying TL contain discontinuous events as indicated by the red boxes. After applying TL, the continuity and smoothness have been improved, as shown in Fig. 16d and 16e. T-2.5D still achieves comparable performance to that of T-3D, with an even higher SSIM value. In summary, T-2.5D has great potential for effectively interpolating randomly missing traces in 3D seismic data from different survey areas, exhibiting great generalization ability.
@@ 4.1 Testing in the Regularly Missing Case
All the interpolated results in the previous section are dealing with irregularly missing traces. In fact, T-2.5D can also be leveraged to handle regular gaps to increase the acquisition density. To validate this, we further assesse the proposed T-2.5D on the Kerry dataset with 67% regularly missing traces (removing two out of every three traces) in Fig. 17e. As shown in Fig. 17b–17d, the interpolation result of T-2.5D shows superior performance to T-2D. Furthermore, it has achieved even better numerical results than T-3D does under such a high missing rate. The residual image in Fig. 17h further demonstrates that T-2.5D can effectively recover signals with little leakage in the regularly missing scenario.
@@ 4.2 Comparison of Trainable Parameters
The number of trainable parameters in a network model is also an important factor for DL-based methods. Models with more trainable parameters can represent more complex functions, but may require more memory and computation (Tu et al., 2024). We further compare the trainable parameters of the three networks. Notably, the total trainable parameters of T-2.5D are composed of two parts: the ones from the four Transformer encoders in the first stage, and the ones from the remainder of T-2.5D in the second stage. As shown in Table 5, the trainable parameters surge from 2D to 3D. Although T-2D has the fewest trainable parameters, its interpolation performance is also inferior to the other two networks. Although the total trainable parameters of T-2.5D are only 75.34% of those of T-3D, T-2.5D can achieve performance comparable to T-3D, or even surpassing it in certain cases. This finding further supports the lightweight and efficient nature of T-2.5D.
@@ 5 Conclusion
Acknowledgements.
This work was jointly supported by the Deep Earth Probe and Mineral Resources Exploration–National Science and Technology Major Project under Grant 2025ZD1007603, and National Natural Science Foundation of China under Grant 42574169.6 Data availability statement
The data used in this paper is available at https://wiki.seg.org/wiki/Open_data. The implementation details of the algorithm are subject to pending intellectual property protection. Therefore, the source code cannot be openly distributed at this stage. Researchers interested in technical discussions may contact the authors for further clarification.